# devoted to memory of Alexei Alexandrovich Starobinsky

### v12

@@ VI From de Sitter thermodynamics to de Sitter decay
$\epsilon_{\rm vac}\sim M_{\rm Pl}^{4}\left(\frac{t_{\rm Pl}}{t+t_{0}}\right)^{2}\,,$ (60)
with the reasonable value of the vacuum energy density in the present time:
$\epsilon_{\rm vac}(t=t_{\rm present})\sim\frac{M_{\rm Pl}^{2}}{t_{\rm present}^{2}}\,.$ (61)
@@
Note that in this approach, the analog of the stiff matter comes from the gravitational degrees of freedom. In this sense it has relations to Refs. Cano2024a ; Cano2024b ; DiFilippo2024 and references therein, where the role of the gravitational degrees of freedom is discussed. Since in this approach the stiff matter is on the order of vacuum energy density, $\epsilon_{\rm stiff}\sim\epsilon_{\rm vac}$, these gravitational degrees of freedom can be responsible for dark matter.Capozziello2012

### v15

@@ VII.3 Connection to holographic principle
$\epsilon_{\rm vac}(t)\sim\epsilon_{M}(t)\sim M_{\rm Pl}^{4}\left(\frac{t_{\rm Pl}}{t+t_{0}}\right)^{2}\,,$ (77)
with the reasonable values of the vacuum energy density and the energy density of matter in the present time:
$\epsilon_{\rm vac}(t_{\rm present})\sim\epsilon_{M}(t_{\rm present})\sim\frac{M_{\rm Pl}^{2}}{t_{\rm present}^{2}}\sim 10^{-120}M_{\rm Pl}^{4}\,.$ (78)
@@
Note that in this approach, the analog of the stiff matter comes from the gravitational degrees of freedom. In this sense it has relations to Refs. Cano2024a ; Cano2024b ; DiFilippo2024 and references therein, where the role of the gravitational degrees of freedom is discussed. Since in this approach the stiff matter is on the order of vacuum energy density, $\epsilon_{\rm stiff}\sim\epsilon_{\rm vac}$, these gravitational degrees of freedom can be responsible both for dark energy and for dark matter.Capozziello2012